A clothing shop makes suits and blazers. Three main resources are used: material, rack space, and labor. The shop has developed this linear programming model for determining the number of suits

and blazers to make ( and ) to maximize profits


maximize Z = 100x1 + 150x2 (profit, $)

Subject to

10x1 + 20x2 smaller than or equal to 300 (material, yd.2

)

x1 + x2 smaller than or equal to 20 (rack space)

10x1 + 4x2 smaller than or equal to160 (labor, hr.)

x1, x2 bigger than or equal to 0


1
Expert's answer
2021-03-31T15:40:35-0400

"\begin{cases}
 10x_1+20x_2\leq 300, \\
 x_1+x_2\leq 20,\\
10x_1+4x_2\leq 160;
\end{cases}"

"Z=100x_1+150x_2\to max,"


"\begin{cases}
 x_1+x_2\leq 20, \\
 10x_1+20x_2\leq 300;
\end{cases}"

"\begin{cases}
 x_1=10, \\
 x_2=10;
\end{cases}"

"Z=100\cdot 10+150\cdot 10=2500" $.